OpenStudy (anonymous):

What is the slope of the line 5x – 2y = 10?

OpenStudy (anonymous):

First solve for y

OpenStudy (anonymous):

5x-10=2y 5/2 x-5=y mx+b=y 5/2=m

OpenStudy (anonymous):

the number next to the x will be your slope